OpenStudy (anonymous):

A tour group split into two groups when waiting in line for food at a fast food counter. The first group bought 7 slices of pizza and 5 soft drinks for $28.42. The second group bought 6 slices of pizza and 4 soft drinks for $23.80. How much does one slice of pizza cost? Write your answer without the $ sign.

OpenStudy (anonymous):

Let price of Pizza be X and SoftDrinks be Y. So the questons would be rephrased as the equation: 7X+5Y = 28.42 - Equation 1 6X+4Y = 23.80 - Equation 2 Find the value of X. Multiply Equation 1 with 4 and Equation 2 with 5, So that the Coefficient of Y is same in both equations. Equation 1 * 4 -> 28X+20Y = 113.68 Equation 2 * 5 -> 30X+20Y = 119 Now Subtracting Equation 1 from Equation 2 2X = 5.32, So X = 5.32/2 = 2.66 Hence the price of slice of Pizza is 2.66